OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (7, -2) and has a slope of -3.

OpenStudy (anonymous):

y - 2 = -3(x + 7) y + 7 = -3(x - 2) y + 2 = -3(x - 7) y - 7 = -3(x + 2)

OpenStudy (btaylor):

\[y-y_1 = m(x-x_1) \rightarrow y+2 = -3(x-7)\] The Third one is your answer.